OpenStudy (anonymous):

a particle is thrown up vertically with a velocity of 50ms.how high would the particle rise and what time would it take to reach the highest point?(g=10ms) particle rise

OpenStudy (anonymous):

formula that you need to use in here for the maximum height is gonna be H=Voy^2/2g.Vo=50ms,theta=90degrees and g=10ms.time here u can calculate by formula t=Voy/g

OpenStudy (anonymous):

\[(v ^{2}-u ^{2})/2g = h\] v = 0, u= 50, g= -10 m/s^2